# : Pixel Grounding Large Multimodal Model

### v2

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models to the vision domain. Initial LMMs used holistic images and text prompts to generate ungrounded textual responses. Recently, region-level LMMs have been used to generate visually grounded responses. However, they are limited to only referring to a single object category at a time, require users to specify the regions, or cannot offer dense pixel-wise object grounding. In this work, we present Grounding LMM (GLaMM), the first model that can generate natural language responses seamlessly intertwined with corresponding object segmentation masks. GLaMM not only grounds objects appearing in the conversations but is flexible enough to accept both textual and optional visual prompts (region of interest) as input. This empowers users to interact with the model at various levels of granularity, both in textual and visual domains. Due to the lack of standard benchmarks for the novel setting of visually Grounded Conversation Generation (GCG), we introduce a comprehensive evaluation protocol with our curated grounded conversations. Our proposed GCG task requires densely grounded concepts in natural scenes at a large-scale. To this end, we propose a densely annotated Grounding-anything Dataset (GranD) using our proposed automated annotation pipeline that encompasses 7.5M unique concepts grounded in a total of 810M regions available with segmentation masks. Besides GCG, GLaMM also performs effectively on several downstream tasks, e.g., referring expression segmentation, image and region-level captioning and vision-language conversations. 11footnotetext: Equally contributing first authors.
@@ 1 Introduction
Detailed region-level understanding requires the laborious process of collecting large-scale annotations for image regions. We propose an automated pipeline to annotate the large-scale Grounding-anything Dataset (GranD) to alleviate the manual labeling effort. Leveraging the automated pipeline with dedicated verification steps, GranD comprises 7.5M unique concepts anchored in 810M regions, each with a segmentation mask. Using state-of-the-art vision and language models, the dataset annotates SAM [18] images through a multi-level hierarchical scheme that enhances annotation quality. With 11M images, 84M referring expressions, and 33M grounded captions, GranD sets a new benchmark in comprehensiveness. In addition to the automatically generated dataset for the GCG, we provide the first high-quality dataset for grounded conversations obtained by revamping the existing manually annotated datasets [37, 16, 49] for GCG using GPT-4 [34] in-context learning. We refer to the high-quality dataset as GranD, denoting its suitability for fine-tuning.
@@
To facilitate model training and evaluation, we create Grounding-anything Dataset (GranD), a large-scale densely annotated dataset. Developed using an automatic annotation pipeline and verification criteria, it encompasses 7.5M unique concepts grounded in 810M regions. Additionally, we propose GranD, a high-quality dataset explicitly designed for the GCG task finetuning, by re-purposing existing open-source datasets.
@@ 3.1 GLaMM Architecture
GLaMM consists of five core components: i) Global Image Encoder, ii) Region Encoder, iii) LLM, iv) Grounding Image Encoder, and v) Pixel Decoder. These components are cohesively designed to handle both textual and optional visual prompts (image level and region), allowing for interaction at multiple levels of granularity and generating grounded text responses (see Fig. 2). These blocks together enable scene-level, region-level, and pixel-level grounding, as explained next. Training specifics are detailed in Appendix A.2.
@@ 4 Data Annotation Pipeline
We introduce our automated annotation pipeline used to create the Grounding-anything Dataset (GranD). GranD is a comprehensive, multi-purpose image-text dataset offering a range of contextual information, from fine-grained to high-level details. It aims to overcome challenges in image understanding and dense pixel-level grounding, thereby expanding the capabilities of visual instruction tuning in LMMs.
@@ 4.2 Relationships and Landmarks (Level-2)
In level-2, multiple short textual descriptions of the overall scene are generated. Phrases extracted from these descriptions are grounded to specific objects in level-1 to form relationships. These relationships articulate connections between multiple objects or define an object’s role within the scene. Further, each scene is assigned a landmark category that includes a primary and a more specific sub-category (see Tab. 7 in Appendix 7).
@@ 4.4 Extra Contextual Insights (Level-4)
Utilizing our automated annotation pipeline, we annotate a corpus of 11M SAM images [18], which are inherently diverse, high-resolution, and privacy-compliant. The resulting dataset comprises 810M regions, each associated with a segmentation mask, and includes 7.5M unique concepts. Further, the dataset features 84M referring expressions, 22M grounded short captions, and 11M densely grounded captions. To our knowledge, this is the first dataset of this scale generated entirely through an automated annotation pipeline (see Tab. 2 for details and Fig. 15 for dataset sample visualizations).
| Dataset | Images | Regions | Concepts | Tokens | Captions |
@@
| Method | refCOCO | refCOCO+ | refCOCOg | |||||
| val | testA | testB | val | testA | testB | val(U) | test(U) | |
| CRIS [47] | 70.5 | 73.2 | 66.1 | 65.3 | 68.1 | 53.7 | 59.9 | 60.4 |
| LAVT [50] | 72.7 | 75.8 | 68.8 | 62.1 | 68.4 | 55.1 | 61.2 | 62.1 |
| GRES [26] | 73.8 | 76.5 | 70.2 | 66.0 | 71.0 | 57.7 | 65.0 | 66.0 |
| X-Decoder [63] | - | - | - | - | - | - | 64.6 | - |
| SEEM [64] | - | - | - | - | - | - | 65.7 | - |
| LISA-7B [21] | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 |
@@
4.5 Building GranD for GCG
Motivated by the need for higher-quality data in fine-tuning stage, we introduce GranD. It contains 214K image-grounded text pairs with 2.5K validation and 5K test samples. GranD comprises two primary components: one subset is manually annotated, and the other subset is derived by re-purposing existing open-source datasets.
@@ 4.5 Building GranD for GCG
In addition, we contribute a minor, high-quality manual annotation set to benchmark the GCG task. Using GranD’s automatic annotations as a base, annotators refine referring expressions to match SAM GT masks, yielding around 1000 focused samples for training and 1000 for evaluation (refer to Appendix D and Fig. 14 for designed prompts and dataset visualizations).
@@ 5 Experiments
We perform quantitative evaluations of GLaMM on six benchmarks: i) Grounded Conversation Generation (GCG), ii) referring-expression segmentation, iii) region-level captioning, iv) image-level captioning, v) conversational-style question answering and vi) phrase grounding. We present the first four benchmarks next, and the remaining are discussed in Appendix B.
Grounded Conversation Generation (GCG). We pretrain GLaMM on GranD dataset followed by fine-tuning on the GranD dataset. The results are presented in Tab. 3 on both validation and test splits of the GranD dataset (refer to Sec. 3.2 and Sec. 4.5 for details). GLaMM shows improved performance compared to baseline methods. Pretrained models for BuboGPT and Kosmos-2 are sourced from official releases, and LISA is adapted and trained on the GranD dataset for the GCG task. GLaMM denotes the variant trained on GranD dataset excluding the 1000 human-annotated images. Qualitative results are shown in Fig. 3 and supplementary Fig. 7.
@@
Refer to Appendix C for qualitative results on six downstream tasks, as well as conditional image generation.
@@ 6 Conclusion
We introduce GLaMM, the first model capable of generating natural language responses intertwined with object segmentation masks, allowing for enhanced multimodal user interactions. Recognizing the lack of standardized benchmarks for visually grounded conversations, we introduce the novel task of Grounded Conversation Generation and establish a comprehensive evaluation protocol. To facilitate research and model development, we create the Grounding-anything Dataset (GranD), a large-scale, densely annotated dataset with 7.5 million unique concepts grounded in 810 million regions. Our automated annotation pipeline ensures the reliability and scalability of this dataset used for our model. In addition to these contributions, we curated a dataset specifically tailored for the GCG task (GranD) by leveraging existing open-source datasets, establishing a high-quality fine-tuning dataset to develop visually grounded conversations. Our model performs well on downstream tasks besides GCG, including region and image captioning, referring segmentation, and vision-language conversations.
@@ A.3 Finetuning on Downstream Tasks
We finetune GLaMM on multiple downstream tasks including GCG, referring expression segmentation, region-level captioning and image-level captioning. For GCG, we finetune our model on GranD dataset. A batch size of 160 is used and the model is trained for 5K iterations in total. It is worth noting that GranD dataset is a combination of multiple open-source datasets that we repurposed for GCG task using GPT4 [34]. Please refer to Appendix. D for the prompts designed to query GPT4 for constructing GranD dataset, along with the dataset visualizations.
@@ A.4.1 LLM Prompts and In-context Learning
Main category |
Fine Category |
Indoor scene |
Living space, Work space, Public space, Industrial space |
Outdoor scene |
Urban landscape, Rural landscape, Natural landscape |
Transportation scene |
Road, Airport, Train station, Port and harbor |
Sports and recreation scene |
Sporting venue, Recreational area, Gym and fitness center |
@@ C.1 Grounded Conversation Generation (GCG)
Fig. 7 shows qualitative results of GLaMM finetuned on GranD dataset. The model could produce dense captions and provide dense pixel-level groundings of the caption.

### v3

# GLaMM: Pixel Grounding Large Multimodal Model

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models to the vision domain. Initial LMMs used holistic images and text prompts to generate ungrounded textual responses. Recently, region-level LMMs have been used to generate visually grounded responses. However, they are limited to only referring to a single object category at a time, require users to specify the regions, or cannot offer dense pixel-wise object grounding. In this work, we present Grounding LMM (GLaMM), the first model that can generate natural language responses seamlessly intertwined with corresponding object segmentation masks. GLaMM not only grounds objects appearing in the conversations but is flexible enough to accept both textual and optional visual prompts (region of interest) as input. This empowers users to interact with the model at various levels of granularity, both in textual and visual domains. Due to the lack of standard benchmarks for the novel setting of visually Grounded Conversation Generation (GCG), we introduce a comprehensive evaluation protocol with our curated grounded conversations. Our proposed GCG task requires densely grounded concepts in natural scenes at a large-scale. To this end, we propose a densely annotated Grounding-anything Dataset (GranD) using our proposed automated annotation pipeline that encompasses 7.5M unique concepts grounded in a total of 810M regions available with segmentation masks. Besides GCG, GLaMM also performs effectively on several downstream tasks, e.g., referring expression segmentation, image and region-level captioning and vision-language conversations. 11footnotetext: Equal contribution.
@@ 1 Introduction
Detailed region-level understanding requires the laborious process of collecting large-scale annotations for image regions. We propose an automated pipeline to annotate the large-scale Grounding-anything Dataset (GranD) to alleviate the manual labeling effort. Leveraging the automated pipeline with dedicated verification steps, GranD comprises 7.5M unique concepts anchored in 810M regions, each with a segmentation mask. Using state-of-the-art vision and language models, the dataset annotates SAM [18] images through a multi-level hierarchical scheme that enhances annotation quality. With 11M images, 84M referring expressions, and 33M grounded captions, GranD sets a new benchmark in comprehensiveness. In addition to the automatically generated dataset for the GCG, we provide the first high-quality dataset for grounded conversations obtained by revamping the existing manually annotated datasets [37, 16, 49] for GCG using GPT-4 [34] in-context learning. We refer to the high-quality dataset as GranDf, denoting its suitability for fine-tuning.
@@
To facilitate model training and evaluation, we create Grounding-anything Dataset (GranD), a large-scale densely annotated dataset. Developed using an automatic annotation pipeline and verification criteria, it encompasses 7.5M unique concepts grounded in 810M regions. Additionally, we propose GranDf, a high-quality dataset explicitly designed for the GCG task finetuning, by re-purposing existing open-source datasets.
@@ 3.1 GLaMM Architecture
GLaMM consists of five core components: i) Global Image Encoder, ii) Region Encoder, iii) LLM, iv) Grounding Image Encoder, and v) Pixel Decoder. These components are cohesively designed to handle both textual and optional visual prompts (image level and region), allowing for interaction at multiple levels of granularity and generating grounded text responses (Fig. 2). These blocks together enable scene-level, region-level, and pixel-level grounding, as explained next. Refer Appendix A.2 for training details.
@@ 4 Data Annotation Pipeline
We introduce our automated annotation pipeline used to create the Grounding-anything Dataset (GranD). GranD is a comprehensive, multi-purpose image-text dataset offering a range of contextual information, from fine-grained to high-level details. It aims to overcome challenges in image understanding and dense pixel-level grounding, thereby expanding capabilities of visual instruction tuning in LMMs.
@@ 4.2 Relationships and Landmarks (Level-2)
In level-2, multiple short textual descriptions of the overall scene are generated. Phrases extracted from these descriptions are grounded to specific objects in level-1 to form relationships. These relationships articulate connections between multiple objects or define an object’s role within the scene. Further, each scene is assigned a landmark category that includes a primary and a more specific sub-category (see Tab. 7 in Appendix A.4.1).
@@ 4.4 Extra Contextual Insights (Level-4)
Utilizing our automated annotation pipeline, we annotate a corpus of 11M SAM images [18], which are inherently diverse, high-resolution, and privacy-compliant. The resulting dataset comprises 810M regions, each associated with a segmentation mask, and includes 7.5M unique concepts. Further, the dataset features 84M referring expressions, 22M grounded short captions, and 11M densely grounded captions. To our knowledge, this is the first dataset of this scale generated entirely through an automated annotation pipeline (see Tab. 2 for details and Fig. 15 in Appendix for dataset sample visualizations).
Dataset | Images | Regions | Concepts | Tokens | Captions† |
@@
Method | refCOCO | refCOCO+ | refCOCOg | ||||||||
val | testA | testB | val | testA | testB | val(U) | test(U) | ||||
CRIS [47] | 70.5 | 73.2 | 66.1 | 65.3 | 68.1 | 53.7 | 59.9 | 60.4 | |||
LAVT [50] | 72.7 | 75.8 | 68.8 | 62.1 | 68.4 | 55.1 | 61.2 | 62.1 | |||
GRES [26] | 73.8 | 76.5 | 70.2 | 66.0 | 71.0 | 57.7 | 65.0 | 66.0 | |||
X-Decoder [63] | - | - | - | - | - | - | 64.6 | - | |||
SEEM [64] | - | - | - | - | - | - | 65.7 | - | |||
LISA-7B [21] | 74.9 | 79.1 | 72.3 | 65.1 | 70.8 | 58.1 | 67.9 | 70.6 | |||
@@
4.5 Building GranDf for GCG
Motivated by the need for higher-quality data in fine-tuning stage, we introduce GranDf. It contains 214K image-grounded text pairs with 2.5K validation and 5K test samples. GranDf comprises two primary components: one subset is manually annotated, and the other subset is derived by re-purposing existing open-source datasets.
@@ 4.5 Building GranDf for GCG
In addition, we contribute a minor, high-quality manual annotation set to benchmark the GCG task. Using GranD’s automatic annotations as a base, annotators refine referring expressions to match SAM GT masks, yielding around 1000 focused samples for training and 1000 for evaluation (refer to Appendix D and Fig. 14 in Appendix for designed prompts and dataset visualizations).
@@ 5 Experiments
We perform quantitative evaluations of GLaMM on six benchmarks: i) Grounded Conversation Generation (GCG), ii) referring-expression segmentation, iii) region-level captioning, iv) image-level captioning, v) conversational-style question answering and vi) phrase grounding. We present the first four benchmarks next, and the remaining are discussed in Appendix B. Grounded Conversation Generation (GCG). We pretrain GLaMM on GranD dataset followed by fine-tuning on the GranDf dataset. The results are presented in Tab. 3 on both validation and test splits of the GranDf dataset (refer to Sec. 3.2 and Sec. 4.5 for details). GLaMM shows improved performance compared to baseline methods. Pretrained models for BuboGPT and Kosmos-2 are sourced from official releases, and LISA is adapted and trained on the GranDf dataset for the GCG task. GLaMM denotes the variant trained on GranDf dataset excluding the 1000 human-annotated images. Qualitative results are shown in Fig. 3 and supplementary Fig. 7.
@@ 6 Conclusion
We introduce GLaMM, the first model capable of generating natural language responses intertwined with object segmentation masks, allowing for enhanced multimodal user interactions. Recognizing the lack of standardized benchmarks for visually grounded conversations, we introduce the novel task of Grounded Conversation Generation and establish a comprehensive evaluation protocol. To facilitate research and model development, we create the Grounding-anything Dataset (GranD), a large-scale, densely annotated dataset with 7.5 million unique concepts grounded in 810 million regions. Our automated annotation pipeline ensures the reliability and scalability of this dataset used for our model. In addition to these contributions, we curated a dataset specifically tailored for the GCG task (GranDf) by leveraging existing open-source datasets, establishing a high-quality fine-tuning dataset to develop visually grounded conversations. Our model performs well on downstream tasks besides GCG, including region and image captioning, referring segmentation, and vision-language conversations.
@@ A.3 Finetuning on Downstream Tasks
We finetune GLaMM on multiple downstream tasks including GCG, referring expression segmentation, region-level captioning and image-level captioning. For GCG, we finetune our model on GranDf dataset. A batch size of 160 is used and the model is trained for 5K iterations in total. It is worth noting that GranDf dataset is a combination of multiple open-source datasets that we repurposed for GCG task using GPT4 [34]. Please refer to Appendix. D for the prompts designed to query GPT4 for constructing GranDf dataset, along with the dataset visualizations.
@@ A.4.1 LLM Prompts and In-context Learning
Main category | Fine Category |
Indoor scene | Living space, Work space, Public space, Industrial space |
Outdoor scene | Urban landscape, Rural landscape, Natural landscape |
Transportation scene | Road, Airport, Train station, Port and harbor |
Sports and recreation scene | Sporting venue, Recreational area, Gym and fitness center |
@@ B.1 Phrase Grounding
![[Uncaptioned image]](extracted/5637309/Figures/supp/phrase_grounding.png)
@@ C.1 Grounded Conversation Generation (GCG)
Fig. 7 shows qualitative results of GLaMM finetuned on GranDf dataset. The model could produce dense captions and provide dense pixel-level groundings of the caption.
@@ Appendix F Ethics and Societal Impact
Appendix G Acknowledgement
The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Alvis partially funded by the Swedish Research Council through grant agreement no. 2022-06725, the LUMI supercomputer hosted by CSC (Finland) and the LUMI consortium, and by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre.